OpenStudy (anonymous):

What are the solutions to the equation 4/x^2 +9/x =-2?

OpenStudy (dumbcow):

multiply everything by x^2 --> 4 +9x = -2x^2

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ \rm 4/x^2 +9/x =-2 }\) mutliply everything time x²

OpenStudy (anonymous):

now same steps as per privious questin..use quadratic formula.

OpenStudy (solomonzelman):

you beat me to it, dumbcow:)

OpenStudy (anonymous):

Can you help me with this other problem @SolomonZelman ? please

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ \rm 4/x^2 +9/x =-2 }\) you get \(\normalsize\color{blue}{ \rm 4 +9x =-2x^2 }\) after multiplying everything times x².

OpenStudy (solomonzelman):

add 2x² to both sides, and solve the quadratic. \(\normalsize\color{blue}{ \rm 2x^2+4 +9x =0 }\)

OpenStudy (anonymous):

A baseball player hits the ball into the air. The height h of the baseball can be modeled approximately with the equation h= -0.0004x^2 +0.15x +4, where x is the horizontal distance from home plate. How far away from home plate did the baseball land? @SolomonZelman or @zaibali.qasmi